All elect lie healer has the resistance R = 10ohm.
Answer #1

a) When they are connected in series the equivalent resistance is:

Req-2R

then the electric power consumed is:

V2 1152 Iseries 66 1.25W series-Reg-2(10)

b) When they are connected in parallel the equivalent resistance is:

R_{eq} = \frac{1}{\frac{1}{R}+\frac{1}{R}} = \frac{R}{2}

then the power consumed is:

V2 、 v115 2645W parallel = 264511

c) To get the maximum heat they must be conncted in parallel.